Check my work 1 Provide the missing data in the following table for a distributor of martial arts products: (Round TurnoverCheck my work 2 Required information [The following information applies to the questions displayed below.] Part 1 of 4 FitnesCheck my work 33 Required information [The following information applies to the questions displayed below.] Part 2 of 4 FitneCheck my work 4 Required information [The following information applies to the questions displayed below.] Part 3 of 4 FitnesCheck my work Required information [The following information applies to the questions displayed below.] Part 4 of 4 FitnessCheck my work 6 Selected sales and operating data for three divisions of different structural engineering firms are given asCheck my work 7 Comparative data on three companies in the same service industry are given below: Required: 2. Fill in the mi
Answer #1


Martial Arts Products Alpha Division Turnover Sales / Average Operating Assets 5 = X/425000 X 2125000 Net Operating Income Sa

Fitness Fanatics Requirement 1 Margin Net Operating Income /Sales = 16940/770000 = 2.20% Turnover Sales / Average Operating A

ROI Margin x Turnover =2.7% x 8.47 =22.87 % Requirement 3 Margin Net Operating Income /Sales (16940 3080)/ 770000 =2.60% Turn

Three Companies Company A Margin Net Operating Income/Sales = 445200 4452000 =10% Turnover Sales / Average Operating Assets =